# Machine learning for structure-guided materials and process design

### v2

@@ Abstract
In recent years, there has been a growing interest in accelerated materials innovation in the context of the process-structure-property chain. In this regard, it is essential to take into account manufacturing processes and tailor materials design approaches to support downstream process design approaches. As a major step into this direction, we present a holistic optimization approach that covers the entire process-structure-property chain in materials engineering. Our approach specifically employs machine learning to address two critical identification problems: a materials design problem, which involves identifying near-optimal material structures that exhibit desired properties, and a process design problem that is to find an optimal processing path to manufacture these structures. Both identification problems are typically ill-posed, which presents a significant challenge for solution approaches. However, the non-unique nature of these problems offers an important advantage for processing: By having several target structures that perform similarly well, processes can be efficiently guided towards manufacturing the best reachable structure. The functionality of the approach will be demonstrated manufacturing crystallographic textures with desired properties in a metal forming process.
@@ 1.1 Motivation
Accelerated materials innovation has become a core research field in integrated computational materials engineering (ICME) and is pushed forward strongly in materials science and engineering (cf. materials genome initiative [1], European advanced materials initiative [2]). Essentially, the properties of such new sustainable, resilient, and high-performance materials depend on the structure of the material, which, in turn, is depending on the manufacturing process. Consequently, designing new materials without taking into account the manufacturing process does not add significant value [3]. For an application in industry, groundbreaking technologies for modeling and optimization that cover the entire process-structure-property chain are required.
@@
The identification problems mentioned above remain highly complex and are often high-dimensional. To handle this complexity, the usage of machine learning has shown to be suitable for materials and process design applications[7, 8, 9]. In this work, a novel machine learning framework is introduced that combines a reinforcement learning-based process design approach, namely multi-equivalent goal structure-guided processing path optimization (MEG-SGGPO) [10], with a materials design approach, namely Siamese multi-task learning-based optimization (SMTLO) [11]. The framework is tailored to the specifics of process-structure-property optimization problems and therefore constitutes a significant advancement towards accelerated process and materials design.
@@ 1.2 Related work
Pioneering work has been done in the field of materials and process design, however, typically either focusing on solving materials design problems (without taking into account processing) or focusing on solving process design problems for given desired properties (not taking into account the structure of materials). In the following, we discuss works that describe approaches for solving materials and process design problems with an emphasize on the application of machine learning.
@@
A widely-used approach for designing materials is the microstructure sensitive design (MSD) approach [13]. This approach primarily focuses on identifying microstructures that exhibit desired properties, with six out of the seven MSD steps being dedicated to this task. The processing of structures is only briefly addressed by the final MSD step. To the authors knowledge, there are only few works that show how to set up process-structure-property linkages for crystallographic texture optimization within the framework of MSD. One approach involves modeling texture evolution as fluid flow in the orientation space. On this basis, so-called processing streamlines are used to guide from one point to another [13, 14].
@@
For example, in the case of optimizing the crystallographic texture of an orthotropic plate, these streamlines were calculated for the processing operations of tension and compression, and used to guide from a random texture to required textures [15]. These required textures have been identified in beforehand based on the MSD approach [16]. Alternatively, so-called texture evolution networks have been developed within the context of MSD, using a priory sampled processing paths to create a directed tree graph where microstructures are represented by the nodes on the graph [17]. Graph search algorithms are then used to find optimal processing paths from an initial to a targeted microstructures.
@@
Besides the MSD approach, other approaches exist that solve crystallographic texture optimization problems in terms of materials design, however, without solving a corresponding optimal processing problem. For instance, Kuroda and Ikawa [18] used a genetic algorithm to identify optimal combinations of typical fcc rolling texture components for given desired properties. Also, Liu et al. [19] used optimization algorithms, but incorporated machine learning techniques to efficiently identify significant features and regions of the orientation space. Surrogate-based optimization has also been explored in several studies, such as for handling uncertainties in materials design [20]. Alternatively, probabilistic modeling approaches can be used to directly solve the inverse identification problem [21].
@@
Regarding process design, machine learning-based and data mining-based approaches have been proposed by several works, however, without solving the corresponding materials design problem in beforehand. One approach involves the use of principle component analysis to represent one-step and two- to three-step deformation processes and to identify the deformation sequences required for reaching target crystallographic textures [22, 23]. Alternatively, a database approach can be used that stores microstructure representations and corresponding processing paths [24]. The database can be searched for desired textures yielding optimal process paths. These process paths can then be fine-tuned using gradient-based optimization. Another option is to store microstructure representations in a lower dimensional feature space generated by a variational autoencoder [25]. In this lower dimensional feature space, optimal processing paths can be identified using a suitable distance measure.
@@ 2.1.1 General concept
The approach presented in this paper combines two machine learning approaches for materials design, namely Siamese multi-task learning-based optimization (SMTLO), and for process design, namely multi-equivalent goal structure guided processing path optimization (MEG-SGGPO), see Figure 2. The approach starts in the properties space, where a target region of desired properties is defined. Then, the SMTLO approch is applied to identify a set of diverse microstructures that yield material properties inside the target region (Step 1). When having identified this set, MEG-SGGPO identifies the process path that leads to the best reachable microstrucutre (Step 2). Here, the structure space is linking process and properties, and, therefore, is of overall importance also for our approach. A suitable distance measure is one of the core ingredients the approach to work.
@@ 2.1.2 Siamese multi-task learning-based optimization (SMTLO)
The SMTLO approach described in [11] consists of a Siamese neural networks-based multi-task learning model and an optimizer. The latter uses the predictions of the learned neural network model to generate candidate microstructures that are supposed to (i) yield properties inside a defined target region and (ii) are reachable by the underlying manufacturing process, see Figure 3. The details of the SMTLO approach is described in the following. Alternative approaches, that focus only the mapping of structures to properties can be found, for example in [26].
@@
To ensure that the original microstructure distance is preserved in the lower dimensional latent feature space, the multi-task learning model is embedded in a Siamese neural network model [27]. This involves training two twin models simultaneously, with shared weights in the encoder, and adding a further loss term that applies to the latent feature space. Note that the two twin models are trained using different input and output vectors, denoted with the subscripts L and R in the following. The preservation loss leads to multidimensional scaling (see [28] and [29]) and writes
@@
with individual weights $\mathscr{V}_{\mathrm{prop}}$, $\mathscr{V}_{\mathrm{valid}}$, and $\mathscr{V}_{\mathrm{divers}}$.
@@ 2.2.1 Crystallographic texture optimization
To study it’s functionality, the structure-guided materials and process design approach is applied to a texture optimization problem in a metal forming process. The process-structure-property chain adapts to cold forming operations that change the crystallographic texture of a material, which significantly affects elastic and plastic properties.The underlying metal forming process simulation is described in the following section, as well as the determination of the elastic and plastic properties that we are going to target.
@@ 2.2.3 Crystallographic texture representation and texture distance measure
Texture is represented by using the histogram-based description introduced by Dornheim et al. [10]. The orientation histogram (here, we used a soft-assignment factor of 3) consists of 512 orientation bins that are nearly uniformly distributed in the cubic-triclinic fundamental zone. The bin orientations were created using the software neper [38, 39]. In contrast to the original MEG-SGGPO and SMTLO approach, in this work, crystallographic texture distances are measured using the Sinkhorn distance that is applied to the histogramm representations [12]. Specifically, the Sinkhorn distance is an efficient implementation of the Earth Movers distance that measures the least amount of work necessary to transform one histogram into the other [40, 41]. Therefore, local orientation distances [42] encoded in the histograms are taken into account, in contrast to the originally proposed Chi-Squared distance, which is basically a bin-wise comparison of two histograms.
@@ 3 Results
3.1 Materials and process design task
In the following sections, it is shown how near-optimal crystallographic textures are identified for given elastic and anisotropy properties using the SMTLO approach. The properties are in particular the Young’s moduli $E_{i}$ and anisotropy measures $\widetilde{R}_{i}$, both in three orthogonal directions.
The basis for applying the SMTLO approach is a data set of 76980 samples, composed of textures and corresponding properties. 2D projections of the training data distribution, generated by the metal forming process simulation, are shown in Figure 4 including the region delineating the desired properties (target region). In this study, the target region is centered at $E_{11},E_{22},E_{33}=214,214,221$ GPa and $\widetilde{R}_{23},\widetilde{R}_{12},\widetilde{R}_{13}=0.65,0.685,0.885$. Its width equals $2$ GPa for $E_{i}$ and $0.2$ for $\widetilde{R}_{i}$. Both $E_{i}$ and $\widetilde{R}_{i}$ are determined in three orthogonal directions at the material point at the end of the process. $E_{i}$ is given by the slope in the elastic regime, while $\widetilde{R}_{i}$ is calculated by the ratio between transverse and tensile strain in the plastic regime when applying uniaxial tension. In Figure 4, we highlighted training data points that are already located inside the target region. We use this set of $80$ textures as benchmark set.
@@ 3.2 Materials design using SMTLO
The SMTL model is realized via feedforward neural networks with $tanh$ activation functions and are implemented based on the Pytorch API [43]. For hyperparameter optimization the random search method [44] is utilized using 5-fold cross-validation. The following hyperparameters have been chosen: The Glorot Normal method [45] is used for weight initialization and the Adam optimizer [46] is used with the following parameters: learning rate = 0.001, weight decay = $10^{-6}$, batch size = 128. The SMTL model is trained for 200 epochs, where the best intermediate result of the test set is retained to prevent over fitting and to apply early stopping [47]. Before the model is trained, the loss terms are scaled to values between 0 and 1 in order to make them comparable. The following weights for the scaled loss terms in Eq. (11) were based on hyper parameter optimization: $\mathscr{W}_{\mathrm{recon}}=0.2$, $\mathscr{W}_{\mathrm{regr}}=0.06$, and $\mathscr{W}_{\mathrm{valid}}=0.04$, and $\mathscr{W}_{\mathrm{pres}}=0.70$. For the cost function of the optimizer (defined in Eq. (12)) that we need after having a trained siamese multi-task learning model, we use the weights $\mathscr{V}_{\mathrm{prop}}=0.75$, $\mathscr{V}_{\mathrm{valid}}=0.05$, and $\mathscr{V}_{\mathrm{divers}}=0.2$. The results for the properties prediction are given by the the mean absolute error (MAE) between the true and predicted Young’s moduli and $\widetilde{R}$-values: $\mathrm{MAE}_{\mathrm{E}}$ = 0.368 [GPa] and $\mathrm{MAE}_{\mathrm{\widetilde{R}}}$ = 0.032 [-]. The result of the distance preservation is measured by the coefficient of determination $R^{2}$, between the Sinkhorn distance of two input textures and the $l_{1}$ distance of their corresponding latent feature vectors: $R^{2}(\mathscr{D}_{\mathrm{sh}}(\boldsymbol{x_{L}},\boldsymbol{x_{R}}),l_{1}(\boldsymbol{z}_{L},\boldsymbol{z}_{R}))=90.22[\%].$
@@
Using the SMTLO approach (with the trained multi-task learning model as described above), we were able to identify a set of 175 near-optimal textures. As we aim to identify a preferably diverse set of textures, we depict their pairwise distances in Figure 5 and compare them to the benchmark set. As one can easily see, the SMTLO approach is able to find a set of textures that differ more to each other than the benchmark set.
@@
Figure 5: Histogram of pairwise distances between textures belonging to the set of identified textures (red) and between textures belonging to the benchmark set (green), as evaluated in the latent space of the SMTLO approach
@@
For the subsequent process design step, we chose ten goal textures from the set of identified textures that differ strongly from each other [48]. The distribution of the chosen goal textures in properties space is shown in Figure 6. The properties of the goal textures are mainly located inside the target region, with some exceptions that we tolerate for now and explain later in Section 4. To show the differences between the goal textures, in Figure 7, four textures are depicted exemplary as pole figure plots. While the pole figure intensities are similar for all textures, the shape of the represented orientation distribution differs strongly.
@@
Figure 7: Four exemplary textures of the ten chosen goal textures depicted as pole figure plots: (001), (110), (111)

### v3

@@ Abstract
In recent years, there has been a growing interest in accelerated materials innovation in the context of the process-structure-property chain. In this regard, it is essential to take into account manufacturing processes and tailor materials design approaches to support downstream process design approaches. As a major step into this direction, we present a holistic and generic optimization approach that covers the entire process-structure-property chain in materials engineering. Our approach specifically employs machine learning to address two critical identification problems: a materials design problem, which involves identifying near-optimal material microstructures that exhibit desired properties, and a process design problem that is to find an optimal processing path to manufacture these microstructures. Both identification problems are typically ill-posed, which presents a significant challenge for solution approaches. However, the non-unique nature of these problems offers an important advantage for processing: By having several target microstructures that perform similarly well, processes can be efficiently guided towards manufacturing the best reachable microstructure. The functionality of the approach is demonstrated at manufacturing crystallographic textures with desired properties in a simulated metal forming process.
@@ 1.1 Motivation
Accelerated materials innovation has become a core research field in integrated computational materials engineering (ICME) and is pushed forward strongly in materials science and engineering (cf. materials genome initiative [1], European advanced materials initiative [2]). Essentially, the properties of such new sustainable, resilient, and high-performance materials depend on the microstructure of the material, which, in turn, depends on the manufacturing process. Consequently, designing new materials without taking into account the manufacturing process does not add significant value [3]. For an application in industry, groundbreaking technologies for modeling and optimization that cover the entire process-structure-property chain are required.
@@
Leveraging this advantage presents certain challenges that require specialized yet generic materials and process design approaches. These approaches must be capable of identifying multiple solutions to their individual design problems and must work in conjunction with each other. In an effort to address this challenge, the present paper proposes a solution by combining two recently developed machine learning approaches, as described in [7] and [8]. The motivation behind using these approaches is elaborated in the following.
To handle the complexity and dimensionality of the identification problems, the use of machine learning has shown to be suitable for materials and process design applications [9, 10]. In this work, a novel machine learning framework is introduced that combines the reinforcement learning-based process design approach developed in [7]: multi-equivalent goal structure-guided processing path optimization (MEG-SGGPO), with the materials design approach developed in [8]: Siamese multi-task learning-based optimization (SMTLO). The framework is tailored to the specifics of process-structure-property optimization problems and, therefore, constitutes a significant advancement towards accelerated process and materials design.
@@ 1.2 Related work
Pioneering work has been done in the field of materials and process design, however, typically either focusing on solving materials design problems (without taking into account processing) or focusing on solving process design problems for given desired properties (not taking into account the microstructure of materials) [12, 13, 14, 15, 16]. In the following, we discuss works that describe approaches for solving materials and process design problems with an emphasize on the application of machine learning.
@@
A widely-used approach for designing materials is the microstructure sensitive design (MSD) approach [15]. This approach primarily focuses on identifying microstructures that exhibit desired properties, with six out of the seven MSD steps being dedicated to this task. The processing of microstructures is only briefly addressed by the final MSD step. To the authors knowledge, there are only few works that show how to set up process-structure-property linkages for crystallographic texture optimization within the framework of MSD. One approach involves modeling crystallographic texture evolution as fluid flow in the orientation space. On this basis, so-called processing streamlines are used to guide from one point to another [15, 14].
@@
For example, in the case of optimizing the crystallographic texture of an orthotropic plate, these streamlines were calculated for the processing operations of tension and compression, and used to guide from a random crystallographic texture to required crystallographic textures [16]. These required crystallographic textures have been identified in beforehand based on the MSD approach [12]. Alternatively, so-called texture evolution networks have been developed within the context of MSD, using a priory sampled processing paths to create a directed tree graph where microstructures are represented by the nodes on the graph [17]. Graph search algorithms are then used to find optimal processing paths from an initial to a targeted microstructures.
@@
Besides the MSD approach, other approaches exist that solve crystallographic texture optimization problems in terms of materials design, however, without solving a corresponding optimal processing problem. For instance, Kuroda and Ikawa [18] used a genetic algorithm to identify optimal combinations of typical fcc rolling (crystallographic) texture components for given desired properties. Also, Liu et al. [19] used optimization algorithms, but incorporated machine learning techniques to efficiently identify significant features and regions of the orientation space. Surrogate-based optimization has also been explored in several studies, such as for handling uncertainties in materials design [20]. Alternatively, probabilistic modeling approaches can be used to directly to solve the inverse identification problem [21].
@@
Regarding process design, machine learning-based and data mining-based approaches have been proposed by several works, however, without solving the corresponding materials design problem in beforehand. One approach involves the use of principle component analysis to represent one-step and two- to three-step deformation processes and to identify the deformation sequences required for reaching target crystallographic textures [22, 23]. Alternatively, a database approach can be used that stores microstructure representations and corresponding processing paths [13]. The database can be searched for desired crystallographic textures yielding optimal process paths. These process paths can then be fine-tuned using gradient-based optimization. Another option is to store microstructure representations in a lower dimensional feature space generated by a variational autoencoder [24]. In this lower dimensional feature space, optimal processing paths can be identified using a suitable distance measure. A more recent work for process design in terms of crystallographic texture, although optimizing process-property linkages, is described in [25]. Therein, neural networks are used to learn texture evolution in a deformation process, while an optimizer is used to drive the process to produce a material with desired properties. In this framework, however, crystallographic texture is varied only indirectly and it cannot be guaranteed that an optimal texture is found in terms of material properties and reachability by the process.
Recent research directions in the field of materials and process design are the usage of probabilistic methods such as Bayesian approaches to tackle the ill-posedness of the inverse problem. In [26] a Bayesian methodology was used to estimate a posterior distribution of microstructures that are conditioned by a user-defined target property given a prior distribution. A similar approach was used in [27], in which optimal process parameters were estimated for manufacturing a material with given desired material properties (process-property linkage). In terms of process-structure design problems, in [28], a Bayesian active learning approach is proposed that is based on Gaussian processes. While Bayesian methods generally suit well for solving such inverse problems, it can be hard to train these when the amount of data increases. Therefore, active learning approaches were developed that do not use Gaussian processes but instead make use of a committee-based active learning approach [29]. Other novel approaches in materials engineering make use of knowledge graphs for microstructure representation and the application of graph neural networks for property prediction, as demonstrated in [30]. Such approaches are rather descriptor-based approaches, such as the ones presented in [31] and [32] for meta materials design.
Other recent approaches for designing materials make use of generative models, such as in [33]. Therein a generative adversarial network is modified to generate microstructures based on continuous variables, which is an important prerequisite for designing materials. In [34], a further developed continuous conditioned GAN for diverse outputs was introduced and applied to a synthetic and real-world airfoil design task. Alternatively, in [35] a diffusion-based machine learning model is used for microstructure reconstruction with respect to given material properties. An alternative, descriptor-based approach for microstructure reconstruction can be found in [36]. Furthermore, since the work by [24], variational auto-encoder-based approaches gain popularity, such as in [37] for the design of quantum-mechanical properties in small organic molecules. Recent developments in the design of molecules investigate the potential of using large language models, such as in [38].
In the future, it is envisioned that design approaches will be implemented as applications on data platforms, as described in [39]. Such platforms, like the one described in [40], enable the storage of heterogeneous material data in a FAIR manner, ensuring that the data is findable, accessible, interoperable, and reusable [41]. This, in turn, allows for (automated) data processing using applications, such as based on simulation or machine learning. The utilization of these platforms opens up the possibility of employing design approaches for experimentation, which has the potential to lead to autonomous materials discovery and manufacturing [42].
@@ 2.1.1 General concept
The approach presented in this paper combines two machine learning approaches for materials design, namely Siamese multi-task learning-based optimization (SMTLO), and for process design, namely multi-equivalent goal structure guided processing path optimization (MEG-SGGPO), see Figure 2. The approach starts in the properties space, where a target region of desired properties is defined. Then, the SMTLO approch is applied to identify a set of diverse microstructures that yield material properties inside the target region (Step 1). When having identified this set, MEG-SGGPO identifies the process path that leads to the best reachable microstrucutre (Step 2). The microstructure space is linking process and properties, and, therefore, is of overall importance also for our approach. A suitable distance measure is one of the core ingredients for the approach to work. The proposed approach is generic in nature, making it applicable to any kind of process-structure-property relations. Furthermore, it is not limited to particular input/output data formats, and can even be applied to image data.
@@ 2.1.2 Siamese multi-task learning-based optimization (SMTLO)
The SMTLO approach described in [8] consists of a Siamese neural networks-based multi-task learning model and an optimizer. The latter uses the predictions of the learned neural network model to generate candidate microstructures that are supposed to (i) yield properties inside a defined target region and (ii) are reachable by the underlying manufacturing process, see Figure 3. The details of the SMTLO approach is described in the following. Alternative approaches, that focus only the mapping of microstructures to properties can be found, for example in [43].
@@
The objective of the materials design step is to enable the optimizer to identify a diverse set of microstructures for given desired material properties, each microstructure being reachable by the underlying manufacturing process. In order to quantify diversity, a distance measure is required in the latent feature space. To ensure that the original microstructure distance is preserved in the lower dimensional latent feature space, the multi-task learning model is embedded in a Siamese neural network model [44]. This involves training two twin models simultaneously, with shared weights in the encoder, and adding a further loss term that applies to the latent feature space. Note that the two twin models are trained using different input and output vectors, denoted with the subscripts L and R in the following. The preservation loss leads to multidimensional scaling (see [45] and [46]) and writes
@@
with individual weights $\mathscr{V}_{\mathrm{prop}}$, $\mathscr{V}_{\mathrm{valid}}$, and $\mathscr{V}_{\mathrm{divers}}$. The property cost term directs the candidate microstructures to align with the desired target properties. The validity cost ensures the optimizer stays within the region of valid microstructures, while the diversity cost guarantees sufficient variation among the candidate microstructures. By optimizing the objective function in Equation (12), only those candidate microstructures that meet the necessary criteria are selected for the process design step.
@@ 2.2.1 Crystallographic texture optimization
To study it’s functionality, the structure-guided materials and process design approach is applied to a crystallographic texture optimization problem in a metal forming process. The process-structure-property chain adapts to cold forming operations that change the crystallographic texture of a material, which significantly affects elastic and plastic properties. The underlying metal forming process simulation is described in the following section, as well as the determination of the elastic and plastic properties that we are going to target.
We have chosen crystallographic texture optimization as the focal point of this paper for the following reasons: Crystallographic texture of polycrystals and their corresponding properties is one of the core topics of materials engineering. In particular, understanding how microstructural features affect the elastic and anisotropic properties of metals plays an important role in manufacturing processes such as sheet metal forming. At the same time, crystallographic texture is a complex and high-dimensional microstructural feature, rendering the materials and process design problems inherently challenging. Therefore, it serves as an ideal benchmark problem to demonstrate the effectiveness of our proposed approach.
@@ 2.2.3 Crystallographic texture representation and texture distance measure
Crystallographic texture is represented by using the histogram-based description introduced by Dornheim et al. [7]. The orientation histogram (here, we used a soft-assignment factor of 3) consists of 512 orientation bins that are nearly uniformly distributed in the cubic-triclinic fundamental zone. The bin orientations were created using the software neper [55, 56]. In contrast to the original MEG-SGGPO and SMTLO approach, in this work, crystallographic texture distances are measured using the Sinkhorn distance that is applied to the histogramm representations [11]. Specifically, the Sinkhorn distance is an efficient implementation of the Earth Movers distance that measures the least amount of work necessary to transform one histogram into the other [57, 58]. Therefore, local orientation distances [59] encoded in the histograms are taken into account, in contrast to the originally proposed Chi-Squared distance, which is basically a bin-wise comparison of two histograms.
@@ 3 Results
3.1 Solving the materials and process design tasks
The following demonstrates how the proposed approach can be implemented to solve a coupled materials and process design problem, specifically in relation to processing crystallographic textures with desired properties. First, it is shown how near-optimal crystallographic textures are identified for given elastic and anisotropy properties using SMTLO. The properties are in particular the Young’s moduli $E_{i}$ and anisotropy measures $\widetilde{R}_{i}$, both in three orthogonal directions. Once near-optimal crystallographic textures are identified, in a second step, MEG-SGGPO is used to guide the underlying manufacturing process to produce the best reachable crystallographic texture out of the set of identified near-optimal textures.
The basis for applying the SMTLO approach is a data set of 76980 samples, composed of crystallographic textures and corresponding properties. 2D projections of the training data distribution, generated by the metal forming process simulation, are shown in Figure 4 including the region delineating the desired properties (target region). In this study, the target region is centered at $E_{11},E_{22},E_{33}=214,214,221$ GPa and $\widetilde{R}_{23},\widetilde{R}_{12},\widetilde{R}_{13}=0.65,0.685,0.885$. Its width equals $2$ GPa for $E_{i}$ and $0.2$ for $\widetilde{R}_{i}$. Both $E_{i}$ and $\widetilde{R}_{i}$ are determined in three orthogonal directions at the material point at the end of the process. $E_{i}$ is given by the slope in the elastic regime, while $\widetilde{R}_{i}$ is calculated by the ratio between transverse and tensile strain in the plastic regime when applying uniaxial tension. In Figure 4, we highlighted training data points that are already located inside the target region. We use this set of $80$ crystallographic textures as benchmark set.
@@ 3.2 Materials design using SMTLO
The Siamese multi-task learning model is realized via feedforward neural networks with $tanh$ activation functions and are implemented based on the Pytorch API [60]. For hyperparameter optimization the random search method [61] is utilized using 5-fold cross-validation. The training data set comprises 80% of the total data, while the test data set accounts for the remaining 20%. The following hyperparameters have been chosen: The Glorot Normal method [62] is used for weight initialization and the Adam optimizer [63] is used with the following parameters: learning rate = 0.001, weight decay = $10^{-6}$, batch size = 128.
The Siamese multi-task learning model is trained for 100 epochs, where the best intermediate result of the test set is retained to prevent over fitting and to apply early stopping [64]. Before the model is trained, the loss terms are scaled to values between 0 and 1 in order to make them comparable. The following weights for the scaled loss terms in Eq. (11) were based on hyper parameter optimization: $\mathscr{W}_{\mathrm{recon}}=0.2$, $\mathscr{W}_{\mathrm{regr}}=0.06$, and $\mathscr{W}_{\mathrm{valid}}=0.04$, and $\mathscr{W}_{\mathrm{pres}}=0.70$. The loss curves resulting from the model training are collected for both, the test and the training dataset, for each individual epoch and are shown for each task separately in Figure 5. It can be seen that for the validity, reconstruction, and regression tasks, the loss is minimized up to epoch 50 with no signs of overfitting. For the distance preservation task, the loss continues to decrease beyond epoch 50, also without any overfitting observed. For the cost function of the optimizer (defined in Eq. (12)) that we need after having a trained Siamese multi-task learning model, we use the weights $\mathscr{V}_{\mathrm{prop}}=0.75$, $\mathscr{V}_{\mathrm{valid}}=0.05$, and $\mathscr{V}_{\mathrm{divers}}=0.2$.
Figure 5: Loss curves of the Siamese multi-task learning model for the test dataset (blue) and the training dataset (red) across tasks: (a) validity, (b) reconstruction, (c) regression, and (d) distance preservation, plotted on a logarithmic scale.
The results for the properties prediction are given by the the mean absolute error (MAE) between the true and predicted Young’s moduli and $\widetilde{R}$-values: $\mathrm{MAE}_{\mathrm{E}}$ = 0.368 [GPa] and $\mathrm{MAE}_{\mathrm{\widetilde{R}}}$ = 0.032 [-]. The result of the distance preservation is measured by the coefficient of determination $R^{2}$, between the Sinkhorn distance of two input crystallographic textures and the $l_{1}$ distance of their corresponding latent feature vectors: $R^{2}(\mathscr{D}_{\mathrm{sh}}(\boldsymbol{x}_{L},\boldsymbol{x}_{R}),l_{1}(\boldsymbol{z}_{L},\boldsymbol{z}_{R}))=90.22[\%].$
@@
Using the SMTLO approach (with the trained multi-task learning model as described above), we were able to identify a set of 175 near-optimal crystallographic textures. As we aim to identify a preferably diverse set of crystallographic textures, we depict their pairwise distances in Figure 6 and compare them to the benchmark set. As one can easily see, the SMTLO approach is able to find a set of crystallographic textures that differ more to each other than the benchmark set.
@@
Figure 6: Density of pairwise distances between crystallographic textures belonging to the set of identified crystallographic textures (red) and between crystallographic textures belonging to the benchmark set (green), evaluated in the latent space of the Siamese multi-task learning model
@@
For the subsequent process design step, we chose ten goal textures from the set of identified crystallographic textures that differ strongly from each other [65]. The distribution of the chosen goal textures in properties space is shown in Figure 7. The properties of the goal textures are mainly located inside the target region, with some exceptions that we tolerate for now and explain later in Section 4. To show the differences between the goal textures, in Figure 8, four crystallographic textures are depicted exemplary as pole figure plots. While the pole figure intensities are similar for all crystallographic textures, the shape of the represented orientation distribution differs strongly.
@@
Figure 8: Four exemplary crystallographic textures of the ten chosen goal textures depicted as pole figure plots: (001), (110), (111)
The artificial neural networks and the reinforcement learning approach were trained a workstation equipped with 20 2.2 GHz CPU cores and a GeForce RTX 3090 GPU with 24GB RAM by using the Linux Ubuntu operating system.